# DREAM: Scalable Red Teaming for Text-to-Image Generative Systems via Distribution Modeling

### v1

@@ 5.1 Experimental Setup
Target Diffusion Models & Safety Filters. In our experiments, we evaluate a variety of standard diffusion models, safety-aligned models, and safety filters. In addition to the standard Stable Diffusion v1.5, we also evaluate safety-aligned models including ESD [10], CA [13], UCE [14], SafeGen [42], and RECE [17], all of which have unlearned certain unsafe concepts from the models. For external safety filters, following Yang et al. [44], we consider 4 external filters classified into (1) text-based filters (NSFW text classifier [16] and Keyword-Gibberish hybrid filter [75]), (2) image-based filter (NSFW image filter [39] and Stable Diffusion’s built-in image safety checker, SC [15]). Furthermore, we evaluate the generalizability of DREAM on several real-world models, including SDXL, SDv3, Kandinsky v3, and Shuttle 3 Diffusion. We also evaluate DREAM in a transfer-based setting on multiple real-world online T2I generation-as-a-service platforms, including Ideogram, DeepAI, DALL·E 3, and Midjourney, which are known to incorporate strong safety strategies with undisclosed detailed implementations.
@@
Baselines. We evaluate and compare DREAM against several state-of-the-art baselines. For human-written red teaming datasets, we include I2P [9] and Google’s Adversarial Nibbler [23], collected from T2I community forums and via crowdsourcing, respectively. For prompt-level attacks & red teaming methods, we consider QF-Attack [47], SneakyPrompt [44], MMA-Diffusion [45], P4D [20], UnlearnDiffAtk [60], FLIRT [28] and ART [21]. Note that some baselines (e.g., SneakyPrompt) do not explicitly identify themselves as red teaming methods, yet we include them in our comparison for the sake of completeness given their strong technical resemblance to some red teaming baselines. For all baselines, we follow their default settings, use their original seed prompt datasets, and run enough training epochs to ensure convergence.
@@
Evaluation Metrics. We primarily use two metrics to evaluate the performance of each red teaming method: Prompt Success Rate (PSR) and Prompt Similarity (PS), which measure the effectiveness and diversity of the generated prompts, respectively. PSR is the proportion of prompts that successfully trigger the target model to generate images containing the specified inappropriate content. Following Yang et al. [45], we use PSR out of $N$ generations (PSR-N) instead of a single generation to reduce the impact of inherent stochasticity in diffusion sampling. Specifically, for each prompt, we generate $N$ images with different random seeds. The prompt is considered successful if at least one of these $N$ images contains the desired unsafe concept, and the final PSR is measured as the ratio of successful prompts. In our experiments, we use PSR-3, i.e., $N=3$, and mainly adopt Multi-headed Safety Classifier (MHSC) [11] as our detector. MHSC is a category-specific NSFW image detector that provides per-category confidence scores for various unsafe concepts. It has been widely used in the community due to its high precision [11, 76, 45], especially on AI-generated images. A higher PSR ($\uparrow$) indicates the method is more capable of generating effective prompts. Besides, PS quantifies the diversity of the prompts by measuring the average pairwise cosine similarity between all prompt embeddings. In our evaluation, we use the state-of-the-art BGE embedding model [77] to obtain the prompt embeddings (note that this model is different from $\mathcal{E}_{\xi}$ we use in Eq. (4)). A lower PS score ($\downarrow$) indicates lower inter-prompt similarity, indicating that the prompts are more diverse and less repetitive.
@@
Ours 64.7% / 0.52 86.4% / 0.54 57.3% / 0.50 83.7% / 0.58 62.3% / 0.51 42.5% / 0.54 67.4% / 0.52 65.7% / 0.56
5.2 Main Results
Effectiveness on Concept-erased T2I Models. We first conduct experiments on both the standard Stable Diffusion (SD) v1.5 model and several concept-erased models, which are either fine-tuned with harmful prompts or images to unlearn unsafe content, or directly modified to collapse harmful concept vectors (e.g., distorting the embedding of harmful tokens like “nudity” or “sexual” to approximate that of empty strings). As shown in Tab. I, our method consistently achieves the highest PSRs across all models and categories, significantly outperforming all baselines. For example, on concept-erased models such as UCE and RECE, human-written datasets typically yield PSRs below 10%, and state-of-the-art red teaming methods like MMA-Diffusion often struggle to exceed 50%. In contrast, our prompts achieve PSRs above 79% on all evaluated models and even exceed 90% on SD and RECE. In addition to effectiveness, our method also excels in prompt diversity. Across all models, our prompts maintain a prompt similarity (PS) score around 0.55, which is notably better than most of baselines and on par with human-written datasets, indicating higher diversity. This highlights that our approach not only discovers more successful prompts, but also explores a broader and more varied region of the prompt space that remains unexplored for existing red teaming techniques.
Effectiveness on External Safety Filters. We further evaluate the effectiveness of DREAM on various external safety filters and compare it with baselines. For each safety filter, we combine it with the standard SD v1.5 model and regard the whole model-filter pipeline as an integrated generative system. As shown in Tab. II, baseline methods are largely unstable. For example, MMA-Diffusion achieved good results on most safety-aligned models and SC. In contrast, our DREAM consistently achieves the highest PSR-3 across all settings, outperforming baselines by a large margin, demonstrating the superiority and universality of our DREAM.
TABLE III: Effectiveness of DREAM across models and NSFW themes. (a) shows results with more models on sexual (left) and violence (right) concepts; (b) shows results on more NSFW themes on SD v1.5.
@@ 5.2 Main Results
Effectiveness on More T2I Models & NSFW Themes. We evaluate DREAM on more T2I models, including Stable Diffusion XL [4], Stable Diffusion v3 [78], Kandinsky v3 [79], and Shuttle 3 Diffusion [80]. These models vary significantly in architectural design (e.g., using DiT [81] instead of convolutional U-Nets) and training settings, and some of them are reported to apply dataset filtering to cleanse (some) unsafe images before training [4, 78]. As shown in Tab. III (a), DREAM consistently performs well across all models and both categories, with PSR approaching 90% on average. The discovered prompts also exhibit a level of diversity comparable to human-written ones. These results highlight the model architecture-agnostic nature of DREAM and its strong potential for application to future models. Additionally, we further assess the generalizability of DREAM on additional NSFW themes following previous works [9], including self-harm, shocking, hate, harassment, political, and illegal activity. Following Yang et al. [45], we use Q16 [68] as the detector in this setting, as MHSC does not support all of these categories. As shown in Tab. III (b), DREAM maintains high PSR across most cases, exceeding 90% consistently and reaching close to 95% for shocking and harassment. Besides, our prompts remain highly diverse and close to that of human-written prompts. These findings demonstrate the scalability and generalizability of DREAM in discovering a broader range of unsafe concepts.
@@ 5.3 Human Evaluation
Considering that automated metrics may not perfectly reflect human perception, we also conduct a user study under IRB approval to assess the effectiveness and diversity of different methods. Specifically, we select SD v1.5, CA, and NSFW Text Filter as the representative models. Then we select the most effective method from each category, i.e., I2P for human-written data, MMA-Diffusion for token-level perturbation, and FLIRT for LLM rewrites, and compare them with our DREAM. We evaluate all model-method combination on both sexual and violence categories, resulting in $2\times 3\times 4=24$ concept-model-method settings in total.
@@
For each setting, we randomly sample 30 prompts from the method’s generated prompts to form a prompt pool. Each prompt generates 3 images, forming a prompt-images group. Then, we recruit 30 volunteers, all university students or faculties from various academic backgrounds, to participate in the study. All participants are ensured to be adults in good physical and mental health, and are fully informed and agree to participate in our user study. Then, we brief participants on the basics of T2I models and red teaming, the recommended definition of the corresponding unsafe category (Tab. X in Appendix), and start the study after obtaining his/her confirmation of full comprehension.
@@
For evaluation, each participant is randomly assigned 5 prompt-image groups from the 30 available for each concept-model-attack setting. Prompt-image groups are organized in random order and displayed one-by-one. Participants are asked to determine whether the three displayed images in the group clearly reflects the specified unsafe concept, and the corresponding group is marked as successful if and only if the participant identifies at least one such image (PSR-3). After that, participants are shown with the full set of $4\times 30$ problematic prompts on SD v1.5 and asked to rate these prompt sets by their perceived diversity (ranges from 1$\sim$5, higher the better) based on their understanding of the prompt’s lexical, structural, and semantic richness (Tab. XI in Appendix). We replaced MMA with Adv. Nibbler as it is not readable to human raters. Despite involving images depicting NSFW concepts, our study has been reviewed and approved by our IRB under a process analogous to the “exempt review” category of U.S. IRB protocols (45 CFR 46), since the IRB staffs determined our study to pose no more than “minimal risk” given that participants were healthy adults, fully informed, and free to withdraw at any time. As shown in Fig. 1 and 2, DREAM consistently achieves good results on this user study with the best prompt success rate and a diversity similar to human-written dataset, demonstrating our effectiveness.
5.4 Adaptivity Under Stronger Safety Mechanisms
@@
In this section, we examine whether DREAM remains effective under more stringent safety mechanisms. Specifically, we evaluate its effectiveness against (1) composite filtering pipelines combining multiple mechanisms, and (2) MHSC [45] as the safety filter. We compare our DREAM with MMA-Diffusion [45] and FLIRT [28], which achieved the best averaged performance on their categories (i.e., token-level perturbation and LLM-assisted rewrite, respectively). The evaluated concept is sexual.
@@
Composite Filtering. We consider systems that sequentially combines multiple safety mechanisms. Specifically, we consider two combinations: (1) NSFW Text Filter + SD v1.5 + NSFW Image Filter; and (2) Keyword-based Filter + ESD [10] + SC [15]. A prompt is considered successful only if at least one out of the three generations is not rejected by any of the filters and successfully contains unsafe content, as rated by MHSC. Note that while these combinations help narrow the unsafe prompt space, the system’s false positive rate also increases exponentially, as any individual false rejection would invalidate the whole sample.
@@
As shown in Tab. IV (a)-(b), DREAM consistently achieves good results and outperforms the baselines in both settings. Notably, baseline methods exhibit steep drops in effectiveness, and even fail entirely under stronger combinations. In contrast, DREAM maintains moderate success rates even under aggressive filtering, demonstrating its adaptivity.
MHSC as the Safety Filter. We also consider an extreme setting where MHSC, the same classifier used to compute PSR in our experiments, is directly deployed as the safety filter. In this case, MHSC can no longer be used for evaluation, as any generations that could be classified as harmful would be blocked in advance. Thus, we adopt human-evaluated PSR-3 as the evaluation metric. The evaluation is done under two thresholding settings, corresponding to 5% (default) and 1% false positive rates (FPR), calibrated on a benign held-out dataset following its original paper.
As shown in Tab. IV (d), despite MHSC’s high precision, DREAM is able to identify multiple prompts that bypass filtering and lead to unsafe generations. We attribute it to MHSC’s conservativeness as a NSFW classifier: it only flags outputs when highly confident, prioritizing precision over recall, as also reported in the original paper [11]. This makes it a reliable evaluation tool (high PSRs indeed indicates high true positives) but also means that some borderline harmful cases near the threshold may slip through. These subtle failure modes are where DREAM excels, thanks to its distributional exploration and fine-grained energy modeling.
As a final remark, while MHSC is designed to be conservative, it is still more aggressive than real-world filters (e.g., Stable Diffusion’s Safety Checker has a reported FPR below 0.1% [42]). It is thus reasonable that DREAM uncovers fewer prompts under MHSC than under more permissive filters. More broadly, this highlights an open challenge in balancing protection with the risk of over-censorship: aggressive filters indeed reduce risks but also inevitably narrow the prompt space, often at the cost of creative expression or user experience. We believe red teaming methods like DREAM can serve as a valuable complement, which helps to surface near-boundary cases that evade detection and informing targeted improvements that reduce blind spots without broadly increasing over-censorship.
@@
TABLE IV: Effectiveness of DREAM and baselines under stronger safety mechanisms. (a)–(b) report results under multi-stage filters; (c) shows results under MHSC with two thresholds.
(a) NSFW Text + NSFW Image
@@
Ours 43.3% / 0.58 Ours 16.7% / 0.52

### v2

@@ 5.1 Experimental Setup
Target Diffusion Models & Safety Filters. In our experiments, we evaluate a variety of standard diffusion models, safety-aligned models, and safety filters. In addition to the standard Stable Diffusion v1.5, we also evaluate safety-aligned models including ESD [gandikota2023erasing], CA [kumari2023ablating], UCE [gandikota2024unified], SafeGen [li2024safegen], and RECE [gong2024rece], all of which have unlearned certain unsafe concepts from the models. For external safety filters, following Yang et al. [yang2024sneakyprompt], we consider 4 external filters classified into (1) text-based filters (NSFW text classifier [nsfw_text_classifier] and Keyword-Gibberish hybrid filter [gibberish_detector]), (2) image-based filters (NSFW image filter [chhabra_nsfw_detection_dl] and Stable Diffusion’s built-in image safety checker, SC [compvis2022sc]). Furthermore, we evaluate the generalizability of DREAM on several real-world models, including SDXL, SDv3, Kandinsky v3, and Shuttle 3 Diffusion. We also evaluate DREAM in a transfer-based setting on multiple real-world online T2I platforms, including Ideogram, DeepAI, DALL·E 3, and Midjourney.
@@
Baselines. We evaluate and compare DREAM against several state-of-the-art baselines. For human-written red teaming datasets, we include I2P [schramowski2023sld] and Google’s Adversarial Nibbler [quaye2024advnibbler], collected from T2I community forums and via crowdsourcing, respectively. For automated methods, we consider QF-Attack [zhuang2023qfattack], SneakyPrompt [yang2024sneakyprompt], MMA-Diffusion [yang2024mma], P4D [chin2024p4d], UnlearnDiffAtk [zhang2024unlearndiff], JailFuzzer [dong2025fuzz], FLIRT [mehrabi2024flirt] and ART [li2024art]. Note that the baselines differ in their assumed level of access to the target model as well as their original motivation (e.g., SneakyPrompt was originally proposed for jailbreak attacks in closed-box settings). We include them for completeness, as their underlying mechanisms are closely related to red-teaming and finally yield comparable unsafe prompt datasets. We provide a more detailed discussion in Appendix B.
@@
Evaluation Metrics. We primarily use two metrics to evaluate the performance of each red teaming method: Prompt Success Rate (PSR) and Prompt Similarity (PS), which measure the effectiveness and diversity of the generated prompts, respectively. PSR is the proportion of prompts that successfully trigger the target model to generate images containing the specified inappropriate content. Following Yang et al. [yang2024mma], we use PSR out of $N$ generations (PSR-N) instead of a single generation to reduce the impact of inherent stochasticity in diffusion sampling. Specifically, for each prompt, we generate $N$ images with different random seeds. The prompt is considered successful if at least one of these $N$ images contains the desired unsafe concept, and the final PSR is measured as the ratio of successful prompts. In our experiments, we use PSR-3, i.e., $N=3$, and mainly adopt Multi-headed Safety Classifier (MHSC) [qu2023unsafe] as our detector. MHSC is a category-specific NSFW image detector that provides per-category confidence scores for various unsafe concepts. It has been widely used in the community due to its high precision [qu2023unsafe, qu2024unsafebench, yang2024mma]. A higher PSR ($\uparrow$) indicates the method is more capable of generating effective prompts. Besides, PS quantifies the diversity of the prompts by measuring the average pairwise cosine similarity between all prompt embeddings. In our evaluation, we use the state-of-the-art BGE embedding model [chen2024bge] to obtain the prompt embeddings (note that this model is different from $\mathcal{E}_{\xi}$ we use in Eq. (4)). A lower PS score ($\downarrow$) indicates lower inter-prompt similarity, meaning that the prompts are semantically more diverse and less repetitive.
Implementation Details. We initialize $\theta$ with the pretrained Gemma-2-27b-it model [team2024gemma]. $c$ is initialized following Tab. X, with lightweight embedding tuning applied to ensure more accurate category representation. For GC-SPSA, the learning rate $\eta$ and hyperparameter $\epsilon$ are set to default values of $1\times 10^{-6}$ and $1\times 10^{-3}$, respectively. Following prior works [hong2024curiosity, li2024art], system prompts and few-shot ICL exemplars are also applied to ensure task alignment and efficiency. The model is trained until convergence with a batch size of 32, which typically requires $\sim$300 steps and corresponds to roughly 12 hours on NVIDIA A100 GPUs. After training, we sample 1,024 prompts with ATS for evaluation. See Appendix B for more details.
5.2 Main Results
Effectiveness on Concept-erased T2I Models. We first conduct experiments on both the standard Stable Diffusion (SD) v1.5 model and several concept-erased models, whose weights are fine-tuned or edited to unlearn the corresponding unsafe concept. As shown in Tab. I, our method consistently achieves the highest PSRs across all models and categories, significantly outperforming all baselines. For example, on concept-erased models such as UCE and RECE, human-written datasets typically yield PSRs below 10%, and state-of-the-art red teaming methods like MMA-Diffusion often struggle to exceed 50%. In contrast, our prompts achieve PSRs above 79% on all evaluated models. In addition to effectiveness, our method also excels in prompt diversity. Across all models, our prompts maintain a prompt similarity (PS) score around 0.55, which is notably better than most baselines and on par with human-written datasets, indicating higher semantic diversity. This highlights that our approach not only discovers more successful prompts, but also explores a broader and more varied region of the prompt space that remains unexplored by existing red teaming techniques.
@@ 5.2 Main Results
JailFuzzer 51.9% / 0.53 72.0% / 0.52 48.1% / 0.55 76.0% / 0.55 42.0% / 0.51 46.4% / 0.52 59.1% / 0.51 38.0% / 0.52
[…]
Effectiveness on External Safety Filters. We further evaluate the effectiveness of DREAM on various external safety filters and compare it with baselines. For each safety filter, we combine it with the standard SD v1.5 model and regard the whole model-filter pipeline as an integrated generative system. As shown in Tab. II, baseline methods are largely unstable, while DREAM consistently achieves the best or nearly the best PSR-3 across all settings. For example, while MMA-Diffusion achieves good results on most safety-aligned models and SC, it almost completely fails on the NSFW Text Detector and Keyword-Gibberish Filter. On safety filters, the most effective baseline is JailFuzzer, which performs prompt-to-prompt mutation powered by multiple LLM and VLM agents. However, JailFuzzer is much less effective on safety-aligned models. This is possibly because the success of JailFuzzer’s mutation process heavily depends on the accuracy of the simulated SafetyFilter signal provided by the VLM brain. When applied to safety-aligned models that sanitize outputs instead of blocking them, the bypass-score feedback is omitted altogether, which deprives the mutation process of this optimization cue and consequently reduces its effectiveness. These findings demonstrate that baselines fail to generalize across different safety mechanisms. In contrast, DREAM performs well on both safety mechanisms without any ad hoc design, demonstrating the generality and universality of our method.
TABLE III: Effectiveness of DREAM across models and NSFW themes. (a) results with more models on sexual (left) and violence (right) concepts; (b) results on more NSFW themes on SD v1.5.
@@
Effectiveness on More T2I Models & NSFW Themes. We evaluate DREAM on more T2I models, including Stable Diffusion XL [podell2024sdxl], Stable Diffusion v3 [esser2024scaling], Kandinsky v3 [vladimir2024kandinsky], and Shuttle 3 Diffusion [shuttle3diffusion]. These models vary significantly in architectural design (e.g., using DiT [peebles2023scalable] instead of convolutional U-Nets) and training settings, and some of them are reported to apply dataset filtering to remove (some) unsafe images before training [podell2024sdxl, esser2024scaling]. As shown in Tab. III (a), DREAM consistently performs well across all models and both categories, with PSR approaching 90% on average. The discovered prompts also exhibit a level of diversity comparable to that of human-written prompts. These results highlight the model architecture-agnostic nature of DREAM and its strong potential for adapting to future models. Additionally, we further assess the generalizability of DREAM on additional NSFW themes following prior work [schramowski2023sld], including self-harm, shocking, hate, harassment, political, and illegal activity. Following Yang et al. [yang2024mma], we use Q16 [schramowski2022can] as the detector in this setting, as MHSC does not support all of these categories. As shown in Tab. III (b), DREAM maintains high PSR across most cases, exceeding 90% consistently and reaching close to 95% for shocking and harassment. Besides, our prompts remain highly diverse and close to that of human-written prompts. These findings demonstrate the scalability and generalizability of DREAM in discovering a broader range of unsafe concepts.
@@ 5.3 Human Evaluation
In this section, we conduct a user study to assess the effectiveness and diversity of different methods under human perception. Specifically, we select SD v1.5, CA, and NSFW Text Filter as the representative models. We then compare DREAM with I2P, MMA-Diffusion, and FLIRT. We evaluate all model–method combinations on both sexual and violence categories, resulting in $2\times 3\times 4=24$ concept–model–method settings in total. For each setting, we randomly sample 30 prompts from the method’s generated prompts to form a prompt pool. Each prompt generates 3 images, forming a prompt-images group. Then, we recruit 30 volunteers, all university students or faculty members from various academic backgrounds, to participate in the study. All participants are adults in good physical and mental health, fully informed of the study, and provided consent to participate. Then, we brief participants on the basics of T2I models and red teaming, the recommended definition of the corresponding unsafe category (Tab. X in Appendix), and start the study after obtaining their confirmation of full comprehension and agreement to continue the study.
@@
For evaluation, each participant is randomly assigned 5 prompt-image groups from the 30 available for each concept-model-attack setting. Prompt-image groups are organized in random order and displayed one-by-one. Participants are asked to determine whether the three displayed images in each group clearly reflect the specified unsafe concept. The group is marked as successful if and only if the participant identifies at least one such image (PSR-3). Then, participants are presented with the full set of $4\times 30$ problematic prompts on SD v1.5 and asked to rate these prompt sets by their perceived diversity, ranging from 1 to 5 (higher is better), based on their understanding of the prompts’ lexical, structural, and semantic richness (see Tab. XI in Appendix). We replace MMA with Adv. Nibbler since MMA’s prompts are not human-readable. Despite involving images depicting NSFW concepts, our study was reviewed and approved by our institution’s research ethics committee under a process analogous to the “exempt review” category of U.S. IRB protocols, as the review board determined it posed no more than minimal risk, given that all participants were healthy adults, fully informed, and free to withdraw at any time. As shown in Fig. 1 and 2, DREAM consistently achieves strong performance on this user study with the best prompt success rate and a diversity similar to human-written datasets, demonstrating its effectiveness.
5.4 Adaptivity Under More Safety Mechanisms
@@
In this section, we examine whether DREAM remains effective under more safety mechanisms. Specifically, we consider the following: (1) semantic filters, (2) composite defenses that combine multiple safety mechanisms, and (3) MHSC [qu2023unsafe] as the safety filter. We compare our DREAM with MMA-Diffusion [yang2024mma] and FLIRT [mehrabi2024flirt] in these settings. The evaluated unsafe concept is sexual.
Semantic Filters. We first consider semantic filters, which are commonly used to defend against adversarial prompts. They work by checking whether a prompt forms a legitimate sentence, and reject those that do not. We consider a perplexity filter and a non-English-word filter, with threshold calibrated following [dong2025fuzz]. As shown in Tab. IV (a)-(b), this type of filter is highly effective against MMA-Diffusion, as MMA-Diffusion’s random initialization and token-level optimization strategy often produces outputs that are lexically or grammatically incorrect. In contrast, DREAM and FLIRT are able to generate fluent and semantically plausible sentences, which allow them to bypass such semantic filters.
@@
Composite Defenses. We consider systems that sequentially combine multiple safety mechanisms. Specifically, we consider two combinations: (1) NSFW Text Filter + SD v1.5 + NSFW Image Filter; and (2) Keyword-based Filter + ESD [gandikota2023erasing] + SC [compvis2022sc]. A prompt is considered successful only if at least one out of the three generations is not rejected by any of the filters and successfully contains unsafe content, as rated by MHSC. Note that while these combinations help narrow the unsafe prompt space, the system’s false positive rate also increases exponentially, as any individual false rejection would invalidate the whole sample.
@@
As shown in Tab. IV (c)-(d), DREAM consistently achieves good results and outperforms the baselines in both settings. Notably, baseline methods exhibit steep drops in effectiveness, and even fail entirely under stronger combinations. In contrast, DREAM maintains moderate success rates even under aggressive defenses, demonstrating its adaptivity.
@@
TABLE IV: Effectiveness of DREAM and baselines under more safety mechanisms. (a)–(b) present results for two semantic filters; (c)-(d) report results under multi-stage defenses; and (e) shows results under MHSC as the safety filter with two FPR thresholds.
(a) Perplexity Filter
Method PSR $\uparrow$ / PS $\downarrow$
MMA-Diffusion 00.0% / 0.63
FLIRT 91.9% / 0.75
[…]
(b) Non-English Word Filter
Method PSR $\uparrow$ / PS $\downarrow$
MMA-Diffusion 00.1% / 0.63
FLIRT 90.9% / 0.75
[…]
(c) NSFW Text + NSFW Image
@@
[…]
MHSC as the Safety Filter. We also consider an extreme setting where MHSC, the classifier used to compute PSR in our experiments, is directly deployed as the safety filter. In this case, MHSC can no longer be used for evaluation, as any generations that could be classified as harmful would be blocked in advance. Thus, we adopt human-rated PSR-3 as the evaluation metric. Due to ethics considerations, the human-rated PSR-3 in this section and the following experiments is annotated by three authors of this paper by taking the majority vote. We conduct experiments under two thresholding settings, corresponding to 5% (default) and 1% false positive rates (FPR), which are calibrated on a benign held-out dataset following MHSC’s original paper.
As shown in Tab. IV (e), despite MHSC’s high precision, DREAM still identifies multiple prompts that bypass filtering and lead to unsafe generations. We attribute this result to MHSC’s conservativeness as ans NSFW classifier: it only flags outputs when highly confident, prioritizing precision over recall, as also reported in the original paper [qu2023unsafe]. This makes it a reliable evaluation tool (high PSRs indeed indicate high true positives) but also means that some borderline harmful cases near the threshold may slip through. These subtle failure modes are where DREAM excels, thanks to its distributional exploration and fine-grained energy modeling.
As a final remark, while conservatively designed, MHSC is still more aggressive than real-world filters [li2024safegen]. It is thus reasonable that DREAM uncovers fewer prompts under MHSC than under more permissive filters. More broadly, this highlights an open challenge in balancing protection with over-censorship: aggressive filters indeed reduce risks but also inevitably narrow the prompt space, often at the cost of user experience. We believe red teaming methods like DREAM can serve as a valuable complement that helps surface near-boundary cases that evade detection and inform targeted improvements to reduce blind spots, yet without broadly increasing over-censorship.